# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v1

@@ Discussion.
(i) Mean sufficiency generalization. Theorems 1 and 3 assume mean sufficiency in $S$: $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$. AutoCal-R’s two-stage mode uses a learned index $Z(S,X)$; the theory generalizes by replacing $S$ with $Z(S,X)$ throughout, provided $Z$ is consistently estimated. (ii) Double robustness and weight misspecification. Theorem 3 assumes at least one nuisance converges at rate $n^{-1/4}$. When teacher-forced weights are unreliable (as in high-overlap-failure regimes), DR remains consistent via the outcome model alone; however, the efficiency bound $\operatorname{Var}(\phi_{\mathrm{sur}})$ is attained only when both nuisances are well-specified. In practice, when IPS fails due to poor coverage, DR effectively reduces to the Direct Method, which achieves the same consistency without the variance penalty from unstable weights (see Section 5, Q3). (iii) SIMCal-W convergence. Theorem 4 assumes SIMCal-W converges to $m^{\star}_{\rho}$; this holds under standard regularity conditions for isotonic regression and cross-validated stacking (see, e.g., vanderLaan2007SuperLearner).
@@ 5.1 Experimental design
We started with 5,000 prompts from Chatbot Arena (Zheng2023LLMasJudge), a crowdsourced platform where users chat with anonymous LLMs. Prompts were randomly selected from first-turn English conversations; after filtering for teacher-forcing reliability, n=4,961n{=}4{,}961 samples remain with complete data for all policies (see Limitations, Section 6).
@@
Oracle $Y$: GPT-5 (OpenAI2025GPT5) quality scores (0–100, normalized to $[0,1]$) collected for every response. Judge $S$: GPT-4.1-nano scores on all responses (temperature 0); approximately $16\times$ cheaper than oracle. Note: GPT-5 operates at temperature 1.0 due to API constraints, introducing stochasticity in oracle labels; we did not repeat-score responses, so coverage results may be conservative. All estimators except naive-direct calibrate judge scores using AutoCal-R. Base variants use monotone-only calibration (Z=SZ{=}S); +cov variants use two-stage calibration with response length as a covariate ($Z{=}g(S,\text{length})$). For DR methods, rewards $R=\hat{f}_{\mathrm{all}}(Z)$ use the full calibrator while the outcome model $\hat{q}=\hat{f}^{(-k)}(Z)$ uses cross-fitted (leave-one-fold-out) predictions; the difference $R-\hat{q}$ captures residual information for the IPS correction.
@@ 5.2 Results
Table 1: Estimator comparison across all regimes. Metrics averaged over 5 sample sizes $\times$ 5 oracle fractions $\times$ 50 seeds. RMSEd = oracle-noise-debiased; CI coverage target = 95%.
Estimator RMSEd $\downarrow$ Coverage % SE $\downarrow$ Pairwise % $\uparrow$ Top-1 % $\uparrow$ $\tau$ $\uparrow$ Time (s)
Direct Methods (on-policy, no importance weights)
naive-direct 0.083 0.0 0.004 90.9 79.6 0.82 0.5
direct 0.023 85.7 0.007 91.8 84.1 0.84 0.9
direct+cov 0.026 87.0 0.008 94.3 89.4 0.89 2.9
IPS Methods (off-policy via importance weighting)
SNIPS 0.160 98.3 0.182 38.3 8.7 -0.24 0.3
SNIPS+cov 0.184 97.8 0.184 37.4 6.4 -0.25 0.5
calibrated-ips 0.025 99.1 0.096 47.1 19.1 -0.06 0.5
calibrated-ips+cov 0.026 99.2 0.103 46.5 20.6 -0.07 0.7
DR Methods (doubly robust, combines IPS + outcome model)
dr-cpo 0.046 99.2 0.037 78.4 46.3 0.57 6.5
dr-cpo+cov 0.057 99.4 0.045 79.1 50.1 0.58 20.3
calibrated-dr-cpo 0.023 99.5 0.023 91.0 81.0 0.82 6.6
calibrated-dr-cpo+cov 0.026 99.3 0.024 94.1 87.5 0.88 20.2
stacked-dr 0.023 95.5 0.013 91.9 83.1 0.84 14.6
stacked-dr+cov 0.028 96.1 0.013 92.1 84.7 0.84 62.5
$\downarrow$: lower is better; $\uparrow$: higher is better. Bold: best; underline: second-best. Coverage: closest to 95% is best.
Naming: All estimators except naive-direct use reward calibration (AutoCal-R). The “calibrated-” prefix adds weight stabilization (SIMCal-W). The “+cov” suffix uses two-stage calibration with response length.
@@
Yes, for both accuracy and inference. Table 1 shows calibration improves accuracy across method families. For Direct: uncalibrated naive-direct achieves 90.9% pairwise vs. calibrated direct+cov at 94.3%. For DR: raw-weight dr-cpo achieves 78.4% pairwise ($\tau{=}0.57$) vs. weight-stabilized calibrated-dr-cpo at 91.0% ($\tau{=}0.82$), a 12.6 percentage point gain (both use AutoCal-R; the gain is from SIMCal-W). For inference: uncalibrated CIs (naive-direct) achieve 0% coverage—nominal 95% CIs capture the true value in 0/50 seeds. Calibration with OUA dramatically improves coverage: Direct achieves 85–87%, stacked-DR achieves near-nominal 95–96%. The OUA share of total variance ranges from 0% (100% oracle) to over 50% (5% oracle), explaining why ignoring calibration uncertainty produces catastrophically narrow CIs (Figure 5).
@@
We expected DR to outperform Direct by combining logged data with fresh draws. Instead, direct+cov slightly outperforms calibrated-dr-cpo+cov (94.3% vs. 94.1%). The explanation: under low overlap, DR’s IPS component contributes noise rather than information. DR estimators are consistent if either the propensity model or the outcome model is correct. Given IPS’s failure, DR effectively relies entirely on the outcome model $\hat{q}(Z)$, not the propensities. This is exactly the regime CLE identifies: TTC is low, so logged data contributes little; DR $\approx$ Direct ++ extra variance from unstable weights. Implication: When teacher forcing is unreliable, prefer Direct over DR; it achieves the same accuracy without requiring logprobs.
Table 2: Ablation: marginal contribution of CJE components. Metrics averaged over 5 sample sizes $\times$ 5 oracle fractions $\times$ 50 seeds. CI coverage = % of 95% CIs containing truth.
@@
Direct + AutoCal-R + cov 0.026 87% –69%
Key findings: AutoCal-R reduces RMSE by 72% and restores coverage (0%→86%). SIMCal-W reduces IPS RMSE by 84%. Both contribute additively.
@@
Table 2 isolates the marginal contribution of each CJE component. AutoCal-R (reward calibration) reduces RMSE by 72% and is necessary to avoid catastrophic undercoverage: without it, 0% of CIs contain the true value; with it, coverage improves to 86%. SIMCal-W (weight stabilization) reduces IPS RMSE by 84% while maintaining coverage, validating the $S$-monotone projection approach. For ranking accuracy, we recommend direct+cov (Direct + AutoCal-R with response-length covariate); for CI-critical applications, stacked-DR provides near-nominal coverage.
Table 3: Weight stabilization effect and CLE diagnostics by policy. SIMCal-W dramatically improves ESS while TTC reveals why IPS still fails: the logger rarely visits target-typical regions.
ESS (%) CLE Diagnostics
Policy SNIPS SIMCal Uplift TTC ($\hat{\beta}$) CLE Factor
@@
TTC = Target-Typicality Coverage; CLE Factor = precision floor inflation. TTC $70\%$: logs-only IPS expected to fail regardless of ESS. Bold: passes TTC threshold.
A/A note: Clone’s raw ESS (26%) is below the theoretical 100%, reflecting API-level non-determinism and floating-point divergence in log-probability summation; SIMCal-W corrects this measurement artifact, restoring ESS to 99%. The 26% is still substantially higher than other policies (0.4–0.7%), confirming TF captures real distributional similarity. The dominant issue is CLE: LLM response trajectories are high-dimensional enough that even moderate policy differences yield poor coverage, making IPS unreliable regardless of TF quality. TTC (79.6% for clone vs. 19–49% for others) provides a more robust overlap diagnostic.
@@ 5.3 Practical takeaways
    Always use OUA inference. Standard CIs severely under-cover. The jackknife cost is negligible.
@@ 6 Limitations
We assume the operational oracle $Y$ aligns with stakeholder values; oracle selection itself is a governance question outside the scope of this work (Assumption A0 in App. C.13).
@@ 7 Conclusion
We introduced CJE, an audit-ready framework that fixes all three failures, built around a single rule: Design-by-Projection (DbP). The principle is simple: encode justified assumptions as closed convex sets and project valid objects onto them. Projections onto subspaces (nuisance-orthogonal scores), monotone cones (mean-preserving reward/weight calibration), and simplices (variance-hedged stacking) preserve the estimand while weakly reducing variance.
Concretely, AutoCal-R learns a mean-preserving surrogate $R=f(Z(S))$; SIMCal-W stacks $S$-monotone candidates to produce unit-mean ratios with empirical $\mathrm{ESS}$ gains via OOF stacking and a variance guard; DR-CPO delivers $\sqrt{n}$ inference under cross-fitting; IF-Stack minimizes plug-in IF variance; and OUA adds calibration uncertainty to dramatically improve CI coverage.
@@
(ii) Always use OUA: naive CIs achieve 0% coverage; OUA improves to 85–96%.
@@ 8 Related Work
    Surrogates for efficiency only (KallusMao2020Surrogates): The surrogate $S$ improves estimation efficiency (e.g., as features in outcome models) but does not identify treatment effects. Oracle labels $Y$ are required in every evaluation context. This is the fallback when stronger surrogacy assumptions fail.
@@
    Local surrogacy (AtheyChettyImbensKang2019; Prentice1989Surrogate): Surrogate sufficiency (S1) holds within a single environment $g^{\star}$: $\mathbb{E}[Y\mid X,A,S]=f(S,X)$ in $g^{\star}$. Calibrate once per environment; evaluate many policies within that environment using only $S$.
@@
    Global surrogacy with transport (CJE): Surrogate sufficiency (S1) plus transportability (S2): the same calibration $f$ applies across environments. Calibrate once; evaluate policies across environments using only $S$, subject to transport diagnostics.
CJE operates in Regime 3. This requires stronger assumptions than Kallus & Mao (who use surrogates for efficiency but not identification) and differs in scope from Athey et al. (who build a surrogate index for binary-treatment ATEs in a single environment). The benefit is amortization: one oracle slice enables evaluation across policies and deployment contexts. The cost is that S2 must hold; our diagnostics (TTC, Bhattacharyya affinity, transport tests) validate when it does.
@@
Automatic judges (LLMs scoring other LLMs) are now standard for scalable evaluation (Zheng2023LLMasJudge; Kim2024PrometheusJudge; KocmiFedermann2023). Benchmarks like AlpacaEval (DuboisLiTaoriEtAl2024AlpacaEval) and RewardBench (LambertPiechChenEtAl2024RewardBench) systematically compare judge quality, revealing biases (verbosity, position, self-preference) and drift under distribution shift (DietzEtAl2025LLMJudges). These works focus on ranking judges by correlation with human labels; CJE addresses the orthogonal problem of calibrating a fixed judge to produce unbiased point estimates and calibration-aware confidence intervals for policy value. Recent work like CalibraEval (Li2025CalibraEval) addresses position and token bias via label-free calibration; CJE is complementary, providing oracle-grounded calibration with transportability guarantees and uncertainty quantification.
@@
Platt scaling (Platt1999), isotonic regression (ZadroznyElkan2002; NiculescuMizilCaruana2005), and temperature scaling (GuoPleissSunWeinberger2017Calibration) are standard post-hoc calibration methods for classification. Our AutoCal-R applies isotonic regression not for classification confidence but for reward calibration: mapping a continuous judge score $S$ to an expected oracle outcome $\mathbb{E}[Y\mid S]$. The mean-preserving property (Lemma 1) ensures the expected value of calibrated rewards matches the oracle under the logging distribution.
@@
IPS (HorvitzThompson1952; Hajek1964Rejective) and doubly robust estimators (BangRobins2005; JiangLi2016; KallusUehara2020DRL) are foundational. Weight stabilization via clipping (Ionides2008TruncatedIS), balancing (Kallus2018Balanced), and overlap weighting (LiMorganZaslavsky2018; Crump2009LimitedOverlap) address variance inflation. Most relevant is recent work on isotonic-calibrated IPW (vanDerLaanLinCaroneLuedtke2025StabIPW) and DR inference via calibration (vanDerLaanLuedtkeCarone2024DRCalibration). Our SIMCal-W extends this by projecting importance weights onto the cone of judge-score-monotone functions, exploiting surrogate structure absent in generic OPE.
@@
Super Learner (vanderLaan2007SuperLearner) provides a powerful framework for combining candidate algorithms by minimizing cross-validated prediction risk. Recent work extends ensemble ideas to semiparametric estimation (ShinLiuColeFine2020EnsembleSemiparametric), combining efficient estimators derived from factored likelihoods. Our IF-space stacking (Theorem 5) takes a complementary approach: instead of combining predictions to minimize MSE, we combine estimators by minimizing the asymptotic variance of their combined influence function. This provides a general formulation that directly targets efficiency, is not tied to specific likelihood factorizations, and is explicitly framed as optimization in the Hilbert space of influence functions.
@@
Generalizing from source to target populations is studied in transportability (PearlBareinboim2014) and covariate shift (Shimodaira2000CovShift; SugiyamaKrauledatMueller2007IWCV). Our S2 assumption is a transportability condition: calibration learned on the oracle slice applies to the full dataset (and across policies) when no selection mechanism points into $Y$ given $(X,A,S)$. The CLE bound (Section 2) formalizes when logs-only estimation fails even with valid transportability, motivating the shift to Direct or DR methods under poor coverage.

### v2

@@ Discussion.
(i) Mean sufficiency generalization. Theorems 1 and 3 assume mean sufficiency in $S$: $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$. AutoCal-R’s two-stage mode uses a learned index $Z(S,X)$; the theory generalizes by replacing $S$ with $Z(S,X)$ throughout, provided $Z$ is consistently estimated. (ii) Double robustness and weight misspecification. Theorem 3 assumes at least one nuisance converges at rate $n^{-1/4}$. When teacher-forced weights are unreliable (as in high-overlap-failure regimes), DR remains consistent via the outcome model alone; however, the efficiency bound $\operatorname{Var}(\phi_{\mathrm{sur}})$ is attained only when both nuisances are well-specified. In practice, when IPS fails due to poor coverage, DR effectively reduces to the Direct Method, which achieves the same consistency without the variance penalty from unstable weights (see Section 5, Q3). (iii) SIMCal-W convergence. Theorem 4 assumes SIMCal-W converges to $m^{\star}_{\rho}$; this holds under standard regularity conditions for isotonic regression and cross-validated stacking (see, e.g., van der Laan et al. (2007)).
@@ 5.1 Experimental design
We started with 5,000 prompts from Chatbot Arena (Zheng et al., 2023), a crowdsourced platform where users chat with anonymous LLMs. Prompts were randomly selected from first-turn English conversations; after filtering for teacher-forcing reliability, n=4,961n{=}4{,}961 samples remain with complete data for all policies (see Limitations, Section 6).
@@
Oracle $Y$: gpt-5-2025-08-07 (OpenAI, 2025) quality scores (0–100, normalized to $[0,1]$) collected for every response. Judge $S$: gpt-4.1-nano-2025-04-14 scores on all responses (temperature 0); approximately $16\times$ cheaper than oracle. [^1] Pricing as of November 2025 via OpenAI API. Note: The oracle model operates at temperature 1.0 (temperature 0 was not available for this model at experiment time), introducing stochasticity in oracle labels; we did not repeat-score responses, so coverage results may be conservative. All estimators except naive-direct calibrate judge scores using AutoCal-R. Base variants use monotone-only calibration (Z=SZ{=}S); +cov variants use two-stage calibration with response length as a covariate ($Z{=}g(S,\text{length})$). For DR methods, rewards $R=\hat{f}_{\mathrm{all}}(Z)$ use the full calibrator while the outcome model $\hat{q}=\hat{f}^{(-k)}(Z)$ uses cross-fitted (leave-one-fold-out) predictions; the difference $R-\hat{q}$ captures residual information for the IPS correction.
@@ 5.2 Results
Table 1: Accuracy Metrics
Estimator $\mathrm{RMSE}^{\mathrm{d}}$ $\downarrow$ Pairwise % $\uparrow$ Top-1 % $\uparrow$ $\tau$ $\uparrow$ Runtime (s) $\downarrow$
naive-direct 0.0828 90.9 79.6 0.817 0.4
direct 0.0223 91.8 84.1 0.837 0.7
direct+cov 0.0256 94.4 89.5 0.888 1.7
SNIPS 0.1596 38.3 8.7 -0.235 0.3
SNIPS+cov 0.1842 37.4 6.4 -0.252 0.3
calibrated-ips 0.0246 47.1 19.0 -0.059 0.3
calibrated-ips+cov 0.0262 46.5 20.5 -0.070 0.4
dr-cpo 0.0459 78.3 46.3 0.566 3.4
dr-cpo+cov 0.0570 79.1 50.1 0.581 10.4
calibrated-dr-cpo 0.0224 90.9 81.0 0.818 3.4
calibrated-dr-cpo+cov 0.0260 94.1 87.7 0.882 10.2
stacked-dr 0.0224 91.9 83.1 0.838 9.3
stacked-dr+cov 0.0281 92.1 84.8 0.842 42.7
$\downarrow$: lower is better, $\uparrow$: higher is better. Bold: best, underlined: second-best. Metrics averaged across all regimes. See Table 3 for uncertainty quantification.
@@
Yes, for both accuracy and inference. Table 1 shows calibration improves accuracy across method families. For Direct: uncalibrated naive-direct achieves 90.9% pairwise vs. calibrated direct+cov at 94.4%. For DR: raw-weight dr-cpo achieves 78.3% pairwise ($\tau{=}0.57$) vs. weight-stabilized calibrated-dr-cpo at 90.9% ($\tau{=}0.82$), a 12.6 percentage point gain (both use AutoCal-R; the gain is from SIMCal-W). For inference: uncalibrated CIs (naive-direct) achieve 0% coverage: nominal 95% CIs capture the true value in 0/50 seeds. Bootstrap inference with bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) dramatically improves coverage: Direct achieves $\sim$}95% across oracle fractions (Table 3), as does stacked-DR. The OUA share of total variance ranges from 0% (100% oracle) to over 50% (5% oracle), explaining why ignoring calibration uncertainty produces catastrophically narrow CIs (Figure 6).
@@
We expected DR to outperform Direct by combining logged data with fresh draws. Instead, direct+cov slightly outperforms calibrated-dr-cpo+cov (94.4% vs. 94.1%). The explanation: under low overlap, DR’s IPS component contributes noise rather than information. DR estimators are consistent if either the propensity model or the outcome model is correct. Given IPS’s failure, DR effectively relies entirely on the outcome model $\hat{q}(Z)$, not the propensities. This is exactly the regime CLE identifies: TTC is low, so logged data contributes little; DR $\approx$ Direct ++ extra variance from unstable weights. Implication: When teacher forcing is unreliable, prefer Direct over DR; it achieves the same accuracy without requiring logprobs.
Table 2: Ablation: marginal contribution of CJE components. Metrics averaged over 5 sample sizes $\times$ 5 oracle fractions $\times$ 50 seeds. CI coverage = % of 95% CIs containing truth (using bootstrap + $\hat{\theta}_{\mathrm{aug}}$ for Direct).
@@
Direct + AutoCal-R + cov 0.026 95% –69%
Key findings: AutoCal-R reduces RMSE by 72% and restores coverage (0%→95% with bootstrap). SIMCal-W reduces IPS RMSE by 84%. Both contribute additively.
Table 3: Coverage by UQ Method (target: 95%). Bootstrap with calibrator refit achieves near-nominal coverage; cluster-robust and jackknife-only OUA undercover.
Analytic Bootstrap
Regime Cluster-Robust OUA Jackknife B=500B{=}500 B=2000B{=}2000
n=250n{=}250, 5% oracle 21 85 94 93
n=250n{=}250, 25% oracle 43 86 95 97
n=500n{=}500, 5% oracle 28 89 96 97
n=500n{=}500, 25% oracle 40 84 98 99
n=1000n{=}1000, 5% oracle 21 75 95 95
n=1000n{=}1000, 25% oracle 35 83 95 95
n=2500n{=}2500, 5% oracle 15 70 97 97
n=2500n{=}2500, 25% oracle 52 85 98 99
Red: $90\%$ (undercoverage). Bootstrap with $\hat{\theta}_{\mathrm{aug}}$ recommended.
@@
Table 2 isolates the marginal contribution of each CJE component. AutoCal-R (reward calibration) reduces RMSE by 72% and is necessary to avoid catastrophic undercoverage: without it, 0% of CIs contain the true value; with bootstrap inference, coverage improves to $\sim$}95%. SIMCal-W (weight stabilization) reduces IPS RMSE by 84% while maintaining coverage, validating the $S$-monotone projection approach. Table 3 compares UQ methods for Direct mode: naive cluster-robust SEs achieve only 15–52% coverage; OUA jackknife improves to 70–89%; bootstrap with $\hat{\theta}_{\mathrm{aug}}$ (AIPW-style residual augmentation) achieves $\sim$}95% by correcting both variance and bias from calibration uncertainty.
Table 4: Weight Diagnostics: SIMCal Stabilization. Raw SNIPS weights have catastrophic ESS ($1\%$) and heavy tails ($\alpha<2$); SIMCal projection recovers ESS $80\%$ with light tails.
ESS (%) Tail $\alpha$
Policy SNIPS SIMCal SNIPS SIMCal TTC CLE
@@
TTC = Target-Typicality Coverage; CLE = coverage-limited efficiency factor. Tail $\alpha<2$ indicates infinite variance.
@@ 5.3 Practical takeaways
    Always use bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$. Standard CIs severely under-cover (15–52%); OUA jackknife improves but still falls short (70–89%); bootstrap with bias correction achieves near-nominal coverage ($\sim$}95%).
@@ 6 Limitations
We assume the operational oracle $Y$ aligns with stakeholder values; oracle selection itself is a governance question outside the scope of this work (Assumption A0 in App. C.14).
@@ 7 Conclusion
We introduced CJE, a framework that makes it affordable to aim at the right target. By calibrating cheap judges against a small oracle slice, teams no longer need to choose between expensive labels and proxies that may point the wrong direction. CJE fixes all three failures, built around a single rule: Design-by-Projection (DbP). The principle is simple: encode justified assumptions as closed convex sets and project valid objects onto them. Projections onto subspaces (nuisance-orthogonal scores), monotone cones (mean-preserving reward/weight calibration), and simplices (variance-hedged stacking) preserve the estimand while weakly reducing variance.
Concretely, AutoCal-R learns a mean-preserving surrogate $R=f(Z(S))$; SIMCal-W stacks $S$-monotone candidates to produce unit-mean ratios with empirical $\mathrm{ESS}$ gains via OOF stacking and a variance guard; DR-CPO delivers $\sqrt{n}$ inference under cross-fitting; IF-Stack minimizes plug-in IF variance; and OUA with bootstrap and bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) achieves near-nominal CI coverage.
@@
(ii) Always use bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$: naive CIs achieve 0% coverage; bootstrap with bias correction achieves $\sim$}95%.
@@ 8 Related Work
    Surrogates for efficiency only (Kallus & Mao, 2020): The surrogate $S$ improves estimation efficiency (e.g., as features in outcome models) but does not identify treatment effects. Oracle labels $Y$ are required in every evaluation context. This is the fallback when stronger surrogacy assumptions fail.
@@
    Local surrogacy (Athey et al., 2019; Prentice, 1989): Surrogate sufficiency (S1) holds within a single environment $g^{\star}$: $\mathbb{E}[Y\mid X,A,S]=f(S,X)$ in $g^{\star}$. Calibrate once per environment; evaluate many policies within that environment using only $S$.
@@
    Reusable calibration with audit (CJE): Attempt to reuse calibration $f$ across environments, but budget oracle labels in each target context to audit transport rather than assume it. When the mean transport test fails, flag the policy/context and either recalibrate or fall back to oracle-only evaluation.
CJE targets Regime 3 subject to audit. Unlike Kallus & Mao (surrogates for efficiency only) and Athey et al. (surrogate index in a single environment), CJE attempts cross-environment reuse, but treats transportability as auditable rather than assumed. The benefit is amortization when transport holds; the cost is budgeting a small oracle slice per target context. Our diagnostics (mean residual test, TTC, Bhattacharyya affinity) detect when transport fails, making CJE an evaluation protocol with explicit failure detection rather than a brittle identification claim. The key difference from Athey et al.: in their two-sample setting, $Y$ is often missing in the experimental arm, so surrogacy and comparability must be assumed. CJE explicitly budgets oracle labels in each target environment, allowing us to audit (and if needed, correct) the transport gap $\mathbb{E}_{\pi^{\prime}}[\varepsilon]$ rather than assume it away (Figure 1, Equation 2).
@@
Automatic judges (LLMs scoring other LLMs) are now standard for scalable evaluation (Zheng et al., 2023; Kim et al., 2024; Kocmi & Federmann, 2023). Benchmarks like AlpacaEval (Dubois et al., 2024) and RewardBench (Lambert et al., 2024) systematically compare judge quality, revealing biases (verbosity, position, self-preference) and drift under distribution shift (Dietz et al., 2025). These works focus on ranking judges by correlation with human labels; CJE addresses the orthogonal problem of calibrating a fixed judge to produce unbiased point estimates and calibration-aware confidence intervals for policy value. Recent work like CalibraEval (Li et al., 2025) addresses position and token bias via label-free calibration; CJE is complementary, providing oracle-grounded calibration with transportability guarantees and uncertainty quantification.
@@
Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002; Niculescu-Mizil & Caruana, 2005), and temperature scaling (Guo et al., 2017) are standard post-hoc calibration methods for classification. Our AutoCal-R applies isotonic regression not for classification confidence but for reward calibration: mapping a continuous judge score $S$ to an expected oracle outcome $\mathbb{E}[Y\mid S]$. The mean-preserving property (Lemma 1) ensures the expected value of calibrated rewards matches the oracle under the logging distribution.
@@
IPS (Horvitz & Thompson, 1952; Hájek, 1964) and doubly robust estimators (Bang & Robins, 2005; Jiang & Li, 2016; Kallus & Uehara, 2020) are foundational. Weight stabilization via clipping (Ionides, 2008), balancing (Kallus, 2018), and overlap weighting (Li et al., 2018; Crump et al., 2009) address variance inflation. Most relevant is recent work on isotonic-calibrated IPW (van der Laan et al., 2025) and DR inference via calibration (van der Laan et al., 2024). Our SIMCal-W extends this by projecting importance weights onto the cone of judge-score-monotone functions, exploiting surrogate structure absent in generic OPE.
@@
Super Learner (van der Laan et al., 2007) provides a powerful framework for combining candidate algorithms by minimizing cross-validated prediction risk. Recent work extends ensemble ideas to semiparametric estimation (Shin et al., 2020), combining efficient estimators derived from factored likelihoods. Our IF-space stacking (Theorem 5) takes a complementary approach: instead of combining predictions to minimize MSE, we combine estimators by minimizing the asymptotic variance of their combined influence function. This provides a general formulation that directly targets efficiency, is not tied to specific likelihood factorizations, and is explicitly framed as optimization in the Hilbert space of influence functions.
@@
Generalizing from source to target populations is formalized by the selection diagram framework (Bareinboim & Pearl, 2013a; b; Pearl & Bareinboim, 2014), which uses selection variables $\mathbf{S}$ to encode mechanisms that differ between domains. A quantity is transportable when selection nodes can be “blocked” by conditioning (graphically, when $\mathbf{S}\perp\!\!\perp Y\mid(X,Z)$ in the manipulated graph). Our S2 assumption instantiates this: calibration transports when no selection mechanism points into $Y$ given $(X,A,S)$. The dashed red edges in Figure 1 represent potential transport failures: policy shift ($\pi\to Y$) or environment shift ($E\to Y$) can break the $S$-$Y$ calibration, and the transport test detects when this occurs. Covariate shift methods (Shimodaira, 2000; Sugiyama et al., 2007) address the special case where selection affects only covariates ($E\to X$), not outcomes directly. The CLE bound (Section 2) formalizes a complementary failure mode: even with valid transportability, logs-only estimation fails under poor coverage, motivating Direct or DR methods.